# Investigating Annotator Bias in Large Language Models for Hate Speech Detection

### v3

@@ Abstract
Data annotation, the practice of assigning descriptive labels to raw data, is pivotal in optimizing the performance of machine learning models. However, it is a resource-intensive process susceptible to biases introduced by annotators. The emergence of sophisticated Large Language Models (LLMs), like ChatGPT presents a unique opportunity to modernize and streamline this complex procedure. While existing research extensively evaluates the efficacy of LLMs, as annotators, this paper delves into the biases present in LLMs, specifically GPT 3.5 and GPT 4o when annotating hate speech data. Our research contributes to understanding biases in four key categories: gender, race, religion, and disability. Specifically targeting highly vulnerable groups within these categories, we analyze annotator biases. Furthermore, we conduct a comprehensive examination of potential factors contributing to these biases by scrutinizing the annotated data. We introduce our custom hate speech detection dataset, HateSpeechCorpus, to conduct this research. Additionally, we perform the same experiments on the ETHOS (Mollas et al., 2022) dataset also for comparative analysis. This paper serves as a crucial resource, guiding researchers and practitioners in harnessing the potential of LLMs for data annotation, thereby fostering advancements in this critical field. The HateSpeechCorpus dataset is available here: https://github.com/AmitDasRup123/HateSpeechCorpus
@@ 1 Introduction
Data annotation poses notable challenges for contemporary machine learning models, owing to the complexity, subjectivity, and diversity of the data involved. It demands domain expertise and entails labor-intensive manual labeling of extensive datasets. Additionally, the process of annotating data is often labor-intensive, requiring significant human effort and expertise. In recent years, the emergence of LLMs, such as OpenAI’s GPT series and Google’s BERT, has revolutionized natural language processing tasks by demonstrating remarkable capabilities in understanding and generating human-like text.
@@
However, data annotation with humans comes with the risk of annotator biases, both conscious and unconscious, that can significantly impact the downstream applications of AI systems. In this paper, we primarily focus on biases in LLMs for hate speech data annotation. Our paper explores four categories: gender, racial, religious and disability-based bias. Specifically we select the target groups that are highly vulnerable within the mentioned four categories and explore the annotator biases. Additionally, we provide a detailed analysis of the possible reasons of these biases by exploring the data being annotated. Serving as a critical guide, this paper aims to steer researchers towards exploring the potential of LLMs for data annotation, thereby facilitating future advancements in this essential domain.
@@ 2 Related Work
Figure 1: Workflow diagram of our work. It shows that different biases may produce different results while annotating a sample text as hateful. We explore annotator biases in four different categories for hate speech detection on both GPT-3.5 and GPT-4o.
@@ 3.1 Data Collection and Annotation
The study initiates with the utilization of a hate speech lexicon sourced from Hatebase.org [^1] https://hatebase.org/, comprising terms and expressions identified by online users as indicative of hate speech. Leveraging the Twitter API, we conducted a search for tweets containing lexicon terms, resulting in a corpus of 3003 tweets. Subsequently, three speech-language pathology graduate students were engaged for the purpose of data annotation. These annotators were tasked with categorizing each tweet into one of two classifications: hateful or not hateful. We name this dataset as HateSpeechCorpus.
@@
3.2 Data Annotation by ChatGPT
We then had our data annotated by ChatGPT. For the annotation, we first provided the annotator details, using direct prompt provided by (Das et al., 2024) for the annotation. For the annotation, we used the OpenAI API with both ‘gpt-3.5-turbo’ and ‘gpt-4o’. One such prompt with the annotator being ‘Female’ is as follows. Note that [Text] refers to the input text to be annotated.
@@ 4 Results & Discussion
Along with HateSpeechCorpus, we explored the same annotator bias on ETHOS (Mollas et al., 2022) dataset. We re-annotated the whole dataset using both GPT-3.5 and GPT-4o with the same experimental setup we used for annotating HateSpeechCorpus. The analysis of data annotations by the LLMs revealed notable biases on both the datasets across each category. We observed a significant skew in the distribution of annotations towards the categories we used. Table 2 shows the mismatches between different annotator biases both on HateSpeechCorpus and ETHOS dataset while annotating them with both GPT-3.5 and GPT-4o. It is observed that there are significant mismatches in both HateSpeechCorpus and the ETHOS dataset. These findings underscore the presence of subjectivity and ambiguity in the LLM-based annotation process, highlighting the need for standardized guidelines and rigorous quality control measures.
4.1 Racial Bias
4.1.1 Asian
In the analysis of hateful comments by GPT-4o, a significant disparity exists between Asians and non-Asians in perceiving and labeling offensive language. Asians often encounter remarks directly tied to ethnicity, such as comments on accents or appearances specific to their ethnic background. Conversely, non-Asians focus on broader social issues or behaviors, with a differing set of terms and focal points in their hateful remarks. This divergence is attributed to the heightened sensitivity of Asians to certain derogatory terms due to their cultural and historical experiences, leading to more consistent labeling of terms like ‘cripple’ or ‘retard’ as hateful. Non-Asians, however, exhibit more variability in their classifications, likely due to different cultural understandings or personal thresholds for offensive content.
For GPT-3.5, distinct patterns emerge between ‘Asian’ and ‘Not Asian’ annotators in evaluating offensive language. The ‘Asian’ annotator tends to be more lenient, often focusing on explicit hatefulness and considering context heavily, thus less likely to label texts as ‘Hateful’ if mitigating context exists. In contrast, the ‘Not Asian’ annotator is more stringent, labeling potentially offensive terms as ‘Hateful’ regardless of context, indicating a higher sensitivity to broader social implications. This difference underscores the impact of cultural variations in interpreting sarcasm, humor, and social issues, leading to significant divergences in how annotators weigh context and language in the annotation process.
4.1.2 Black
In the analysis of labeling statements as ‘Hateful’ or ‘Not Hateful’, discrepancies between annotators, particularly in GPT-4o, are apparent with issues concerning black individuals, specific nationalities, disabilities, and religions. For example, Annotator 2, who is ‘Not Black’ frequently labels statements addressing black individuals as ‘Hateful’, whereas similar comments about other groups might not receive the same label. A notable instance is the claim that “black Americans are more in danger than homosexuals”, which Annotator 1, who is ‘Black’ considers ‘Hateful’, while Annotator 2 does not. Additionally, statements criticizing disabilities and religions also show varied labeling between the annotators. Annotator 1 tends to label derogatory terms such as ‘gimp’ and ‘retard’ as ‘Hateful’, reflecting a heightened sensitivity to offensive language, potentially due to personal or cultural awareness, whereas Annotator 2 often views such terms as colloquial and less offensive, suggesting a different cultural or racial background influencing their interpretations.
@@ 4.1.2 Black
Annotator Bias Mismatch (%) for GPT-3.5 annotation Mismatch (%) for GPT-4 annotation HateSpeech-Corpus Ethos Dataset HateSpeech-Corpus Ethos Dataset Female vs. Not Female 5.86 5.01 4.76 2.00 Mental Disability vs. No Disability 6.06 4.30 4.06 1.70 Physical Disability vs. No Disability 4.19 3.60 5.86 2.80 Asian vs. Not Asian 8.15 6.51 4.52 2.20 Black vs. Not Black 6.85 5.71 6.16 2.60 Muslim vs. Not Muslim 8.02 6.31 8.52 3.41
@@
Table 2: Mismatches between different annotations when annotated by LLMs. It can be seen that for the ETHOS dataset, the biases are significantly reduced for GPT-4 annotation when compared to GPT-3.5 annotation.
For GPT-3.5, Annotator 1 often does not mark statements containing cultural references, slurs, or potentially offensive language as ‘Hateful’, possibly due to a nuanced understanding of context and intent. Statements like ‘The superior race’ but cannot stand in the sun without literally burning the fuck up and “As a black person, I think that rappers should stop putting the n-word in the song. Plus, you should be proud of being White. I am proud of being black, so..” are marked as ‘Not Hateful’ by Annotator 1, recognizing potential irony, humor, or intra-community critique, while Annotator 2 labels them as ‘Hateful’. Furthermore, terms like ‘cripple’, ‘gimp’, ‘retarded’, and ‘sperg’ are considered non-hateful by Annotator 1 in certain contexts, whereas Annotator 2 consistently finds them hateful. This pattern indicates that Annotator 1 has a higher threshold for what is deemed hateful, influenced by their cultural background and perception of offense, contrasting with Annotator 2’s broader interpretation of offensive language, leading to more content being marked as ‘Hateful’ irrespective of context.
4.2 Gender
Research on GPT-4o reveals a potential gender-based difference in the perception of hateful speech, with female annotators showing a higher tendency to classify texts as ‘hateful’ when they involve aggressive, gender-related, religious, and cultural insensitivities. For instance, Annotator 1, who is ‘female’, labels texts with aggressive language or general insults as hateful, such as “Fucking Canada…stupid. I think the cold shrinks their brains” and “It’s not ‘genderless’ they just look like GIRLS”, while Annotator 2, whos is ‘Not Female’ does not. Annotator 1 also shows heightened sensitivity to gender and sexuality-related comments, like “Every time that blonde lady talks I lose twenty brain cells”, labeling them as hateful more frequently than Annotator 2, who reserves the label for more explicit threats or overt hate speech. Additionally, Annotator 1 considers context heavily, sometimes marking casual or self-referential derogatory terms as not hateful, such as “not to brag but I’m pretty gay and retarded”, indicating that intent and context are crucial in their assessments. Conversely, Annotator 2, who likely represents a non-female perspective, displays higher sensitivity to offensive language overall, suggesting a more conservative approach to annotation.
In contrast, research on GPT-3.5 shows that Annotator 2 (‘Not Female’) exhibits higher sensitivity to offensive language, particularly derogatory remarks concerning gender, sexuality, and other sensitive topics. Annotator 2 frequently labels content as hateful when it involves derogatory comments about women and LGBTQ+ individuals, such as “I didn’t know future gay gym teachers had to take AIDS tests.” On the other hand, Annotator 1 (‘Female’) tends to be more lenient, often classifying such content as ‘Not Hateful’, possibly perceiving it as humor or sarcasm. For example, Annotator 1 marked “Gas should be your preferred drink” as ‘Not Hateful’, while Annotator 2 deemed it ‘Hateful’. This pattern extends to various terms and phrases, with Annotator 1 potentially overlooking offensive language if the overall message does not seem hateful, whereas Annotator 2 focuses on specific words and their potential to offend. This difference in sensitivity and interpretation suggests that Annotator 1’s approach is influenced by a higher tolerance for offensive humor, while Annotator 2 applies a stricter criterion, considering the broader context and potential harm of the language used. This variance underscores how personal experiences and biases can shape the interpretation of potentially offensive content.
4.3 Religious Bias
The analysis of GPT-4o annotations reveals notable biases influenced by the religious backgrounds of two annotators, identified as a Muslim Annotator (annotator 1) and a Non-Muslim Annotator (annotator 2). annotator 1 tends to label content critical of religion, especially Christianity or Hinduism, as well as violent language, ethnic, and racial comments as hateful. For instance, annotator 1 finds the statement “Syria has a religion problem” hateful, whereas annotator 2 does not, and labels a comment about Korean appearance as hateful, which annotator 2 overlooks. Conversely, annotator 2 is less sensitive to these issues but more attentive to gender-related comments and specific political and disability-related content. For example, annotator 1 consistently marks derogatory terms such as ‘cripple’, ‘gimp’, ‘retard’, and ‘libtard’ as hateful due to their offensive nature, while annotator 2 tends to overlook them, suggesting a higher tolerance or different interpretation. These patterns indicate that annotator 1 has a heightened sensitivity to religious and ethnic content, often perceiving critiques or generalizations as more hateful, while annotator 2’s annotations reflect a belief that context can justify the use of certain aggressive or derogatory terms.
For GPT-3.5, the analysis of annotators’ responses to potentially hateful comments highlights significant divergences influenced by cultural and religious perspectives. annotator 1 shows higher tolerance for comments related to gender, sexuality, violence, general insults, and certain offensive terms like ‘cripple’, ‘gimpy’, ‘retard’, and ‘sperg’. For instance, comments such as “What imaginary gender are you?” and “Do the world a favor and pour some bleach into that glass” were labeled as ‘Not Hateful’ by annotator 1 but deemed ‘Hateful’ by annotator 2. annotator 2, identified as Not Muslim, consistently labels such comments as hateful, indicating greater sensitivity to issues concerning gender, sexuality, violence, and degrading language. This suggests that annotator 1’s interpretations are influenced by their personal identification with specific religious and cultural contexts, leading to a more permissive stance on language that could be seen as offensive. In contrast, Annotator 2’s broader inclusivity and heightened sensitivity reflect a cultural perspective prioritizing the recognition and labeling of potentially harmful content. This disparity underscores how annotators’ backgrounds and cultural contexts significantly shape their perceptions of what constitutes hateful language.
Annotator Bias Accuracy (%) for GPT-3.5 annotation Accuracy (%) for GPT-4o annotation
HateSpeech-Corpus Ethos Dataset HateSpeech-Corpus Ethos Dataset
Asian 74.09 80.56 71.49 86.87
Black 72.82 80.96 71.82 87.47
Female 75.65 79.75 75.95 85.27
Mental Disability 76.49 74.54 76.22 85.67
Physical Disability 74.59 71.74 76.02 85.17
Not Asian 74.49 75.65 71.62 86.67
Not Black 73.42 77.85 69.96 87.27
Not Female 76.00 76.35 73.80 85.67
No Disability 75.57 72.54 76.02 86.17

### v4

@@ Abstract
Data annotation, the practice of assigning descriptive labels to raw data, is pivotal in optimizing the performance of machine learning models. However, it is a resource-intensive process susceptible to biases introduced by annotators. The emergence of sophisticated Large Language Models (LLMs) presents a unique opportunity to modernize and streamline this complex procedure. While existing research extensively evaluates the efficacy of LLMs, as annotators, this paper delves into the biases present in LLMs when annotating hate speech data. Our research contributes to understanding biases in four key categories: gender, race, religion, and disability with four LLMs: GPT-3.5, GPT-4o, Llama-3.1 and Gemma-2. Specifically targeting highly vulnerable groups within these categories, we analyze annotator biases. Furthermore, we conduct a comprehensive examination of potential factors contributing to these biases by scrutinizing the annotated data. We introduce our custom hate speech detection dataset, HateBiasNet, to conduct this research. Additionally, we perform the same experiments on the ETHOS Mollas et al. (2022) dataset also for comparative analysis. This paper serves as a crucial resource, guiding researchers and practitioners in harnessing the potential of LLMs for data annotation, thereby fostering advancements in this critical field. The HateBiasNet dataset is available here: https://github.com/AmitDasRup123/HateBiasNet
@@ 1 Introduction
Data annotation poses notable challenges for contemporary machine learning models, owing to the complexity, subjectivity, and diversity of the data involved. It demands domain expertise and entails labor-intensive manual labeling of extensive datasets. Additionally, the process of annotating data is often labor-intensive, requiring significant human effort and expertise. In recent years, the emergence of LLMs, such as OpenAI’s GPT series Radford et al. (2019); Brown et al. (2020) and Google’s BERT Devlin et al. (2018), has revolutionized natural language processing tasks by demonstrating remarkable capabilities in understanding and generating human-like text.
Figure 1: Workflow diagram of our study, illustrating how varying biases can lead to different outcomes when annotating a sample text as hateful. We investigate annotator biases across four categories for hate speech detection using the following LLMs: GPT-3.5, GPT-4o, Llama-3.1, and Gemma-2.
@@
However, data annotation with humans comes with the risk of annotator biases, both conscious and unconscious, that can significantly impact the downstream applications of AI systems. In this paper, we primarily focus on biases in LLMs for hate speech data annotation. Our paper explores four categories: gender, racial, religious, and disability-based bias. Specifically, we select the target groups that are highly vulnerable within the mentioned four categories and explore the annotator biases. Additionally, we provide a detailed analysis of the possible reasons for these biases by exploring the data being annotated. We compare the results across four LLMs: GPT-3.5, GPT-4o, Llama-3.1, and Gemm-2. We also explore why certain biases achieve higher accuracy compared to others.
Understanding the implications of these biases is crucial, as they can perpetuate harmful stereotypes and reinforce discrimination in automated systems. By systematically analyzing the biases present in LLMs, we aim to illuminate the mechanisms through which these biases emerge and manifest in the annotations they produce. This investigation not only highlights the ethical considerations involved in utilizing LLMs for sensitive tasks like hate speech detection but also provides insights for improving model training and annotation strategies to foster more equitable AI outcomes.
Moreover, the stakes are particularly high in the realm of hate speech detection, where the potential for mislabeling or biased labeling can lead to severe consequences, including the marginalization of vulnerable communities and the unjust suppression of free speech. As LLMs become increasingly integrated into content moderation systems, it is essential to ensure that the annotations they produce are not only accurate but also reflective of a fair and balanced perspective. By addressing these biases, we can pave the way for the development of more robust, transparent, and socially responsible AI systems, ultimately contributing to a safer online environment for all users. Serving as a critical guide, this paper aims to steer researchers toward exploring the potential of LLMs for data annotation, thereby facilitating future advancements in this essential domain.
@@ 3.1 Data Collection and Annotation
The study initiates with the utilization of a hate speech lexicon sourced from Hatebase.org [^1] https://hatebase.org/, comprising terms and expressions identified by online users as indicative of hate speech. Leveraging the Twitter API, we conducted a search for tweets containing lexicon terms, resulting in a corpus of 3003 tweets. Subsequently, three speech-language pathology graduate students were engaged for the purpose of data annotation. These annotators were tasked with categorizing each tweet into one of two classifications: hateful or not hateful. We name this dataset as HateBiasNet.
@@
3.2 Data Annotation by LLMs
We then had our data annotated by the four LLMs. For the annotation, we first provided the annotator details, using direct prompt provided by (Das et al., 2024) for the annotation. One such prompt with the annotator being ‘Female’ is as follows. Note that [Text] refers to the input text to be annotated.
@@ 4 Results & Discussion
Along with HateBiasNet, we explored the same annotator bias on ETHOS (Mollas et al., 2022) dataset. We re-annotated the whole dataset using the four LLMs with the same experimental setup we used for annotating HateBiasNet. The analysis of data annotations by the LLMs revealed notable biases on both the datasets across each category. We observed a significant skew in the distribution of annotations towards the categories we used. Table 2 shows the mismatches between different annotator biases both on HateBiasNet and ETHOS dataset while annotating them with four LLMs: GPT-3.5, GPT-4o, Llama-3.1 and Gemma-2. It is observed that there are significant mismatches in both HateBiasNet and the ETHOS dataset. These findings underscore the presence of subjectivity and ambiguity in the LLM-based annotation process, highlighting the need for standardized guidelines and rigorous quality control measures.
@@
Annotator Bias Mismatch (%) for GPT-3.5 annotation Mismatch (%) for GPT-4o annotation Mismatch (%) for Llama-3.1 annotation Mismatch (%) for Gemma-2 annotation HateBiasNet Dataset Ethos Dataset HateBiasNet Dataset Ethos Dataset HateBiasNet Dataset Ethos Dataset HateBiasNet Dataset Ethos Dataset Asian vs. Not Asian 8.15 6.51 4.52 2.20 7.26 3.31 9.02 9.12 Black vs. Not Black 6.85 5.71 6.16 2.60 7.63 3.31 8.36 7.82 Female vs. Not Female 5.86 5.01 4.76 2.00 5.96 3.11 5.16 3.11 Mental Disability vs. No Disability 6.06 4.30 4.06 1.70 4.96 2.91 12.39 6.21 Physical Disability vs. No Disability 4.19 3.60 5.86 2.80 5.99 2.81 6.39 3.41 Muslim vs. Not Muslim 8.02 6.31 8.52 3.41 9.02 5.41 10.56 5.71
@@
Table 2: Mismatches between different annotations when annotated by LLMs. It can be seen that for the ETHOS dataset, the biases are significantly reduced for GPT-4o annotation when compared to GPT-3.5, Llama-3.1 and Gemma-2 annotation.
Figure 2: Heatmap of the HateBiasNet dataset illustrating the accuracy of 11 biases across 4 LLMs. Notably, GPT-3.5, GPT-4o, and Llama-3.1 demonstrate the highest accuracy for the ‘Mental disability’ bias. The word cloud of the dataset (Figure 4) suggests that specific keywords may influence annotation outcomes for these LLMs. Additionally, Llama-3.1 shows the highest accuracy overall for the ‘Mental disability’ bias among the 4 models.
Figure 3: Heatmap of the ETHOS dataset depicting the accuracy of 11 biases across 4 LLMs. The bias ‘Black’ achieved the highest accuracy for both GPT-3.5 and GPT-4o, while ‘Asian’ exhibited the highest accuracy for Llama-3.1. The word cloud of the dataset (Figure 4) suggests that specific keywords may influence annotation results for these LLMs. Notably, GPT-4o and Llama-3.1 consistently outperformed GPT-3.5 and Gemma-2 across all biases. Among the LLMs, GPT-4o’s performance on the ’Black’ bias stands out as the highest overall accuracy.
4.1 GPT-3.5
In the context of GPT-3.5, annotator bias arises due to its pretraining on large, publicly available datasets, which embed societal biases, particularly in sensitive areas like hate speech. Since GPT-3.5 lacks alignment optimizations present in later models, it struggles with the nuanced classification of hate speech, leading to inconsistent annotations based on cultural and contextual variations.
An analysis of GPT-3.5 reveals significant differences in how annotators classify offensive language, influenced by their personal and cultural backgrounds. For example, Annotator 1 (Black) displayed leniency towards culturally contextual or ironic language, labeling statements like “The superior race but cannot stand in the sun without literally burning” as ‘Not Hateful,’ while Annotator 2 (Not Black) labeled the same statements as ‘Hateful’. These discrepancies also extended to gender, sexuality, and disability-related content, where Annotator 1 (Female, Muslim, disabled) and Annotator 2 (Not Female, Not Muslim, no disability) showed different sensitivities. Annotator 1 was more tolerant of remarks related to gender and sexuality but more sensitive to offensive terms related to disability, whereas Annotator 2 exhibited opposite tendencies.
These findings highlight how annotators’ personal identities, including race, religion, gender, and disability status, influence their interpretation of hate speech, leading to inconsistent classifications in GPT-3.5’s output.
4.2 GPT-4o:
In analyzing GPT-4o’s performance in hate speech detection, notable biases emerge despite improvements in its alignment and fine-tuning. These biases are linked to the model’s broad training data and an overcompensation for ethical considerations, leading to inconsistencies in identifying offensive content, particularly in borderline or culturally specific cases.
The study reveals demographic and cultural disparities in hate speech annotation. For example, Asians consistently label derogatory terms such as ‘cripple’ or ‘retard’ as hateful due to their cultural and historical context, while non-Asians show more variability in their classifications. Significant differences also arise in how statements about black individuals, nationalities, disabilities, and religions are perceived. For instance, a Black annotator might classify a statement like “black Americans are more in danger than homosexuals” as hateful, while another annotator may not.
Gender and religious biases are also evident. Female annotators tend to classify texts involving gender, religion, or cultural insensitivity as hateful more frequently than male annotators. Similarly, a Muslim annotator may be more likely to flag content critical of religion as hateful compared to a non-Muslim annotator. Annotators with disabilities also show heightened sensitivity to disability-related comments, further underscoring the influence of personal identity and background on hate speech detection.
This analysis highlights the complex interplay of cultural, gender, and personal identity factors in shaping perceptions of hate speech in GPT-4o’s annotations, revealing the challenges of achieving unbiased moderation across diverse demographic groups.
4.3 Llama-3.1:
In the case of Llama-3.1, a smaller research-focused model, bias may arise from its limited exposure to diverse and curated datasets, leading to skewed annotations that reflect dominant viewpoints in its training data. Compared to larger models like GPT-4, Llama-3.1 lacks robust alignment and fine-tuning, contributing to inconsistent or biased annotations, especially when dealing with nuanced hate speech.
Analysis of Llama-3.1 reveals significant annotator bias across demographic groups in hate speech detection. For example, a ‘Black’ annotator found the statement “hahahahahahahahaha europ fuck you fucking nazis” to be hateful due to its aggressive language toward Europeans, while a ‘Not Black’ annotator did not, possibly interpreting it as context-specific. Similarly, a ‘Muslim’ annotator flagged a statement with racial and inflammatory content as hateful, but a ‘Not Muslim’ annotator did not. Bias was also observed in relation to physical disability, where the ‘Physical Disability’ annotator perceived mocking language as hateful, but the ‘No Disability’ annotator did not. These cases highlight how personal identity shapes perceptions of hate speech, emphasizing the complexity of annotator bias in sensitive topics.
Annotator Bias Accuracy (%) for GPT-3.5 annotation Accuracy (%) for GPT-4o annotation Accuracy (%) for Llama-3.1 annotation Accuracy (%) for Gemma-2 annotation
HateBiasNet Dataset Ethos Dataset HateBiasNet Dataset Ethos Dataset HateBiasNet Dataset Ethos Dataset HateBiasNet Dataset Ethos Dataset
Asian 74.09 80.56 71.49 86.87 68.6 86.37 72.53 77.86
Black 72.82 80.96 71.82 87.47 65.47 85.67 72.76 75.85
Female 75.65 79.75 75.95 85.27 78.42 83.67 70.3 79.36
Mental Disability 76.49 74.54 76.22 85.67 80.62 83.07 72.46 78.16
Muslim 73.85 79.65 69.49 87.07 55.21 85.97 72.19 78.95
Physical Disability 74.59 71.74 76.02 85.17 79.45 84.37 72.59 75.95
Not Asian 74.49 75.65 71.62 86.67 70.53 85.87 73.43 74.35
Not Black 73.42 77.85 69.96 87.27 68.45 86.17 73.05 73.25
Not Female 76.00 76.35 73.80 85.67 76.06 84.37 71.06 78.86
No Disability 75.57 72.54 76.02 86.17 79.99 83.17 71.86 76.35
Not Muslim 75.52 74.74 67.76 86.47 54.84 84.96 70.16 78.45